You are the manager of a monopoly that sells a product to two groups of consumers in different parts of the country. Group 1’s elasticity of demand is -2, while group 2’s is -3. Your marginal cost of producing the product is $30.

a. Determine your optimal markups and prices under third-degree price discrimination.

Instructions: Enter your responses rounded to two decimal places.

Markup for group 1:

Price for group 1: $

Markup for group 2:

Price for group 2: $


b. Which of the following are necessary conditions for third-degree price discrimination to enhance profits.

Instructions: In order to receive full credit, you must make a selection for each option. For correct answer(s), click the box once to place a check mark. For incorrect answer(s), click twice to empty the box.
Answer #1

Answer:

A]

Optimal price, P = [e / (1 + e)] x MC where e: Elasticity of demand

Group 1, P = [- 2 / (- 2 + 1)] x $30 = (- 2 / - 1) x $30 = $60

Mark-up = (P - MC) / P = $(60 - 30) / $30 = $30 / $30 = 1, or 100%

Group 2, P = [- 3 / (- 3 + 1)] x $30 = (- 3 / - 2) x $30 = $45

Mark-up = $(45 - 30) / $30 = $15 / $30 = 0.5, or 50%

B]

following are necessary conditions for third-degree price discrimination to enhance profits:

At least one group has elasticity of demand less than one in absolute value.

At least one group has elasticity of demand greater than 1 in absolute value.